OpenStudy (anonymous):

How do you find a square root? example √ 96 how do you get 4√ 6 ?

OpenStudy (callisto):

\[\sqrt{96} = \sqrt{16\times 6} = \sqrt{16}\times \sqrt{6} = ...\]

OpenStudy (anonymous):

but why 16 and 6? and why is there a 4?

OpenStudy (anonymous):

\[\sqrt{a^{2}*b}==a \sqrt{b}\]

OpenStudy (anonymous):

You need to get the highest square number which in this case is 16. There are no other square numbers higher than 16 that multiplied by another number will get you 96. The four Is there cause it is the square root of 16.

OpenStudy (anonymous):

and then what do you do after you get 16?

OpenStudy (anonymous):

oh, never mind :) thanks

OpenStudy (anonymous):

No worries :)